OpenStudy (anonymous):

WILL GIVE FAN AND MEDAL Find an equation for the nth term of the arithmetic sequence. -15, -6, 3, 12, ...

OpenStudy (anonymous):

We know that the equation is increasing by 9 isn't?

OpenStudy (anonymous):

Yes

alones (alones):

@jewea just gave you a hint that they all go by nine "12-(9)=3; 3-(9)=-6 u see??

OpenStudy (anonymous):

I understand that part

alones (alones):

okayii so then what u don't understand

alones (alones):

just add till you get on the ninth term i'll help u

OpenStudy (anonymous):

an = -15 + 9(n + 1) an = -15 x 9(n - 1) an = -15 + 9(n + 2) an = -15 + 9(n - 1) these are my answer choices and idk how they got this from the question

alones (alones):

if it's fine with you

OpenStudy (anonymous):

Im very confused

alones (alones):

I think the first one cuz it goes 3,12,21,30... the second one doesnt' really match the question it goes 0,-135,-270.. #3_12,21.... #4 is RIGHT :)

OpenStudy (anonymous):

Thankyou :)